OpenStudy (anonymous):

Answer please: 45divided by 9+8*2-5+3

OpenStudy (anonymous):

0.2

OpenStudy (anonymous):

I got 1.9

OpenStudy (anonymous):

9+8*2-5+3=1.95

OpenStudy (anonymous):

No it equals 23

OpenStudy (anonymous):

i was thinking rounding to the nearest tenth

OpenStudy (anonymous):

well i mean the final answer my bad

OpenStudy (anonymous):

you'd multiply 8*2 =16 9+16-5+3=23 45/23=1.95

OpenStudy (anonymous):

ya u were right my bad

OpenStudy (anonymous):

Ha it's ok just don't want her to get her homework wrong

OpenStudy (anonymous):

ya same

OpenStudy (anonymous):

thxs guys

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

yup good luck stina deserves a metal

OpenStudy (anonymous):

just did lol

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

ooo I feel special xD

OpenStudy (anonymous):

thanks